16.        This morning, TL Trucking invested $69,000 to help fund a company expansion project planned for 5 years from now. How much additional money will the firm have 5 years from now if it can earn 9 percent rather than 5 percent on its savings compounded monthly?
Answer #1

We use the formula:
A=P(1+r/12)^12n
where
A=future value
P=present value
r=rate of interest
n=time period.

At 9%:

A=69000*(1+0.09/12)^(12*5)

=69000*1.56568103

=108031.991

At 5%:

A=69000*(1+0.05/12)^(12*5)

=69000*1.28335868

=88551.7489

Hence additional money=108031.991-88551.7489

=$19480.24(Approx)